OpenStudy (anonymous):

In a candy bar, 165 calories are from sugar. There are 250 calories in the candy bar. What percent of the calories are from sugar?

OpenStudy (anonymous):

165/250=.66 66%

OpenStudy (anonymous):

percentage= (amt of sugar)/total calories * 100 =165/250 * 100 =66%

OpenStudy (anonymous):

THANKS!

OpenStudy (across):

This is a covert way of telling you that chocolate will murder you.